Tony invested in two bank accounts, the first with 5% annual interest and the second with 8% annual interest. For the one with 8%, he deposited $1500 more than the one with 5%. After a year his interest was $510. How much did Tony deposit on each account originally?
Answer by richwmiller(17219)

We don't know the total invested.
We know the difference in the two accounts.
We know that the interest for the two accounts is $510
0.08*x+0.05*y=510
We know that the account at 8% has $1500 more.
x=1500+y
We substitute for x
0.08*(1500+y)+0.05*y=510
We multiply out
120+0.08y+0.05*y=510
We combine like terms.
0.13*y=390
Isolate y
y=$3000 at 5%
x=1500+y
Calculate x
x=$4500 at 8%
Now,we know the total invested is: 7500
Total invested $4500+$3000=$7500
We check
0.08*4500+0.05*3000=510
360+150=510
510=510
Since this statement is TRUE and neither x nor y is negative all is well.
codeintmt